OpenStudy (anonymous):

If f(2) = 3 and f'(2)=-1 find an equation of the tangent line when x = 2. How would I start this??

OpenStudy (anonymous):

use the point slope formula \[y-y_1=m(x-x_1)\] with \[x_1=2,y_1=3, m=-1\]

OpenStudy (anonymous):

ahh, thank you. i got y=5-x?

OpenStudy (anonymous):

i didn't do it

OpenStudy (anonymous):

\[y-3=-(x-2)\] \[y-3=-x+2\] \[y=-x+5\] yup

OpenStudy (anonymous):

You need to learn order your equation by variable

OpenStudy (anonymous):

thank youu! your explanation was very helpful!

OpenStudy (anonymous):

the form is y = mx + b

OpenStudy (anonymous):

@Quyen do like that to mention the right person!

OpenStudy (anonymous):

There're so many great solver here, as long as you wish to learn, you'll be better in no time :) Now back to work !

OpenStudy (anonymous):

:) thanks

OpenStudy (anonymous):

You can post whole day whole night here, learning for FREE :)

OpenStudy (anonymous):

it's awesome!

OpenStudy (precal):

I agree with Chlorophyll even though AP calculus will give credit for the other formula. They will even award credit if you just substitute and do not simplify